OpenStudy (candycanegone):

A carpenter purchased 70 ft of redwood and 90 ft of pine for a total cost of $323. A second purchase, at the same prices, included 100 ft of redwood and 50 ft of pine for a total cost of $375. Find the cost per foot of redwood and of pine.

OpenStudy (wolf1728):

70R + 90P =$323 100R + 50P $375 There's your equations. Solve and get your answers.

OpenStudy (candycanegone):

how do i solve?

OpenStudy (candycanegone):

like 170r + 140p ?

OpenStudy (wolf1728):

No- you first have to eliminate an unknown. I messed up - will have to start over

OpenStudy (candycanegone):

it's okay i got it! :)

OpenStudy (candycanegone):

thank you!

OpenStudy (wolf1728):

a) 70R + 90P =$323 b) 100R + 50P $375 Multiplying equation b by -.7 b) -70R -35P = -262.50 Now, add equation a + b a) 70R + 90P =$323 b) -70R -35P = -262.50 55P = 60.50 P = 1.10 The cost of pine

OpenStudy (wolf1728):

and to solve for price of redwood a) 70R + 90*1.10 =$323 70R = 323 - 99 70R = 224 R = 3.20 price of redwood